# TRiSM for Agentic AI: A Review of Trust, Risk, and Security Management in LLM-based Agentic Multi-Agent Systems

### v3

@@ Abstract
Agentic AI systems, built upon large language models (LLMs) and deployed in multi-agent configurations, are redefining intelligence, autonomy, collaboration, and decision-making across enterprise and societal domains. This review presents a structured analysis of Trust, Risk, and Security Management (TRiSM) in the context of LLM-based Agentic Multi-Agent Systems (AMAS). We begin by examining the conceptual foundations of Agentic AI and highlight its architectural distinctions from traditional AI agents. We then adapt and extend the AI TRiSM framework for Agentic AI, structured around four key pillars: Explainability, ModelOps, Security, Privacy and Governance, each contextualized to the challenges of multi-agent LLM systems. A novel risk taxonomy is proposed to capture the unique threats and vulnerabilities of Agentic AI, ranging from coordination failures to prompt-based adversarial manipulation. To support practical assessment in Agentic AI works, we introduce two novel metrics: the Component Synergy Score (CSS), which quantifies the quality of inter-agent collaboration, and the Tool Utilization Efficacy (TUE), which evaluates the efficiency of tool use within agent workflows. We further discuss strategies for improving explainability in Agentic AI , as well as approaches to enhancing security and privacy through encryption, adversarial robustness, and regulatory compliance. The review concludes with a research roadmap for the responsible development and deployment of Agentic AI, outlining critical directions to align emerging systems with TRiSM principles for safe, transparent, and accountable operation.
@@ 1 Introduction
1.1 Background and Motivation
@@
The global market for Artificial Intelligence (AI) agents is projected to grow from \$5.4 billion in 2024 to \$7.6 billion in 2025  [1]. By mid-2025, over 70% of enterprise AI deployments are expected to involve multi-agent or action-based systems, reflecting a dramatic shift from traditional single-agent or rule-based conversational models [2]. An AI agent, typically defined as a computational entity that perceives its environment and takes actions to achieve goals  [3], has undergone a fundamental transformation. While early agents were task-specific and deterministic, modern systems have evolved significantly ,increasing operational autonomy by over 70% on benchmark tasks [4]. These Agentic AI systems now integrate large language models (LLMs), external tools, and persistent memory to support complex planning and coordinated decision-making [5].
This rapid evolution presents new opportunities but also brings challenges in coordination, control, transparency, and risk assessment across heterogeneous agents. In particular, as the shift toward agentic AI architectures amplifies and evolves, longstanding concerns around safety, fairness, interpretability, and systemic harms resurface. While such issues have been central to machine learning (ML) systems broadly, they become more promiment in Agentic AI, where autonomous agents operate in high-stakes domains like healthcare, law, and public services, areas where their decisions can carry profound and lasting societal consequences.
@@
From Isolated Agents to Agentic AI The autonomous software agents date back to the 1990s [6], however, the advent of LLM-driven multi-agent ecosystems [7] since 2023 has made “Agentic AI”, a distinctly new and rapidly expanding research focus [5]. Traditional agents primarily focussed on automating limited tasks, such as information retrieval [8], data summarization [9], and dialogue response [10]. Typically driven by single step logic [11] or scripted rules, these traditional agents lacked the deep reasoning, adaptability, and persistence required for complex, multi-step problem solving .
@@
In contrast, Agentic AI consists of collaborative agents with specialized roles (e.g., planner, coder, analyst), and enabled by LLMs and tool use [12]. These systems dynamically decompose tasks, share context, and pursue high-level goals across long timelines [13]. This shift reflects not only a change in technology but also a fundamental increase in complexity and autonomy, marking the emergence of machine collectives capable of exhibiting emergent and decentralized behavior.
@@ 1.2 Trust, Risk, and Security Challenges in Agentic AI
The rise of Agentic AI introduces new challenges that extend beyond those of traditional ML systems. For example, unlike deterministic agents [14], LLM-based multi-agent systems exhibit emergent and often opaque decision-making, increasing the risk of cascading errors, biases, and unintended behaviors. These characteristics necessitate a robust approach to have trust and ensuring trustworthiness, in high-stakes domains such as healthcare, finance, and law. Below, we discuss some of these challanges:
@@
    Trust and Trustworthiness In the context of Agentic AI, trust refers to a user willingness to rely on an AI system, while trustworthiness denotes whether the system consistently behaves in a safe, fair, and predictable manner, thereby deserving that trust. This distinction is critical in high-stakes settings, where undue reliance on an untrustworthy system can lead to severe consequences, such as misdiagnoses in healthcare or biased rulings in legal applications. The stochastic nature of LLM reasoning [15] further complicates trustworthiness, as randomness in output generation, while promoting creativity and ensemble diversity, introduces inconsistency and non-repeatability. These factors hinder traceability, verification, and compliance auditing, all of which are essential for trustworthy systems.
@@
    Security Risks in Multi-Agent Systems Security risks escalate as autonomous agents gain access to tools, external APIs, and persistent memory, amplifying the potential for privacy breaches, adversarial misuse, or regulatory violations in the absence of robust human oversight [16]. For example, a multi-agent supply chain optimizer may autonomously coordinate between procurement and logistics agents but inadvertently leak sensitive data or violate compliance protocols (e.g., GDPR or HIPAA) if safeguards are inadequate. The distributed and cooperative nature of multi-agent systems introduces unique vulnerabilities that traditional security measures often fail to address.
@@
    Specific Threats: Prompt Injection, Spoofing, and Impersonation LLM-based agents are susceptible to novel security threats exploiting their language-based interfaces and cooperative behaviors. One prominent threat is prompt injection [17], where attackers design inputs with hidden or malicious instructions to manipulate agent behavior. Recent studies [18] describe “prompt infection” as a phenomenon where malicious prompts propagate from one agent to another, akin to a computer virus spreading across networks [19]. Such corrupted outputs can lead to cascading effects, including data leaks, fraudulent transactions, misinformation, or coordinated misbehavior within an agent society.
    Another critical vulnerability involves spoofing and impersonation [20]. Spoofing means faking an entity’s identity or credentials (e.g., using a false name or address) to deceive a system, while impersonation involves actively pretending to be a specific agent or user to gain trust, access, or privileges. In multi-agent systems, where agents rely on credentials or tokens for authentication during coordination, an adversary could steal an API key or mimic a trusted peer to issue unauthorized commands or access sensitive information.
@@
    Privacy Risks in Multi-Agent LLM Systems LLM-based agents often need to handle user data, proprietary business information, or other sensitive inputs to fulfill their tasks [21]. In a multi-agent context, this challenge is amplified by the fact that agents may share information with each other (e.g. via a shared memory store or message passing) to collaborate.
@@
Traditional evaluation and safety frameworks, designed for deterministic agents AI systems, are inadequate for addressing the autonomy and control of Agentic AI. To address these gaps, frameworks like AI TRiSM (Artificial Intelligence Trust, Risk, and Security Management) [22, 23] proposes lifecycle-level controls, including explainability, secure model orchestration, and privacy management. AI TRiSM is a framework to help organizations build AI systems that people can trust. The goal is to focus not just on technical performance, but also on ethical values like fairness, transparency, and security. The framework brings together practices that support governance, strengthen reliability and robustness, and protect data and make sure that AI systems are not only effective, but also aligned with human values.
Figure 1: Taxonomy of TriSM for Agentic AI systems presented in this review. It includes five major sections: Fundamentals of Agentic AI Systems ( Agentic AI vs. Traditional AI Agents, LLM-Based Architecture), The TRISM Framework (Governance, Explainability, ModelOps, Application Security, Privacy), Threats and Risks , Trust and Explainability, and Evaluation Metrics (including Human-Centered and Composite Metrics).
1.3 Problem Statement
Agentic AI systems increasingly rely on the coordination of multiple LLM-powered agents to perform complex, long-horizon tasks. However, no unified framework exists for managing the trust, risk, and security across dynamic as well as multi-agent workflows. Recent accidents, such as, an autonomous research agent citing unverifiable medical sources or collaborative swarms generating inconsistent forecasts due to memory drift, highlight that traditional safeguards are inadequate [24]. These traditional systems are vulnerable to orchestration breakdowns, agent collusion, data leakage, and opaque decision-making pathways. To ensure safety and reliability in Agentic AI systems, a structured TRiSM perspective is therefore urgently needed.
1.4 Scope and Objectives
Objectives This study is a review article that maps the TRiSM (Trust, Risk, and Security Management) [22] framework to LLM-powered, multi-agent “Agentic AI” systems. We structure the discussion around four foundational pillars of TRiSM [25]: explainability, security, lifecycle governance, and privacy. The objective of this review is to offer researchers and practitioners a coherent and actionable reference for the design, deployment, and governance of trustworthy Agentic AI ecosystems.
For the purposes of this study, we define Agentic AI as Multi-Agent Systems (MAS) [26] powered by LLMs that exhibit autonomous planning, tool use, memory retention, and emergent reasoning capabilities, with or without human supervision. These LLM-based agents can dynamically adapt workflows based on interactions with tools, users, and other agents, distinguishing them from traditional symbolic or rule-based agents.
Scope This paper is a conceptual survey and synthesis. It does not propose a new architecture for Agentic AI, nor does it include synthetic benchmark evaluations. Instead, it integrates current perspectives on AI TRiSM to inform future research and the responsible development of Agentic AI systems. The scope of this paper is confined to LLM based Agentic AI systems and AMAS  [^3] Throughout this paper, AMAS designates multi-agent systems built on LLMs.
Necessity of this Survey Agentic AI research has largely centered on agent modeling, planning, and task coordination, while comparatively overlooking critical dimensions such as adversarial robustness, lifecycle governance, and system-level explainability. As these systems are increasingly deployed in high-stakes domains - such as healthcare, scientific research, and finance the absence of an integrated TRiSM framework exposes stakeholders to opaque decision-making and unmanaged risks. This survey addresses that gap by synthesizing TRiSM principles in the context of LLM-driven multi-agent systems and providing actionable guidance for researchers, engineers, and policymakers.
To the best of our knowledge, few existing surveys systematically explore LLM-based multi-agent systems through the lens of trust, risk, and security. Table 1 underscores this gap by comparing prior reviews with respect to their coverage of TRiSM dimensions.
1.5 Contributions
Our key contributions in this review are:
  • We propose the first TRiSM framework for Agentic AI systems. This framework is explicitly tailored to systems in which several LLM-powered agents collaborate and cooperate. The framework is organized around pillars: Explainability, ModelOps, Security, Privacy and Governance around autonomous decision to guide the development and use of safe and ethical Agentic AI systems. (see Fig. 1 and Sec. 6).
  • We offer a systematic taxonomy of risks that emerge when multiple agents interact (Sec. 5). This taxonomy discusses prompt injection, memory poisoning, collusive failure, emergent misbehavior, and related threats commonly seen in AMAS.
  • We introduce an evaluation dimensions for Agentic AI systems that covers Trustworthiness, Explainability, User-centered Performance, Coordination, and Composite Assessment (Sec. 8). To support this template, we define two new metrics: the Component Synergy Score (CSS), which measures the quality of inter-agent collaboration, and the Tool Utilization Efficacy (TUE), which measures how correctly and efficiently agents invoke external tools. Our aim is to achieve high performance in both dimensions, accuracy and usability, without sacrificing one.
  • We contribute to the management and control of Agentic AI by mapping existing TRiSM techniques such as explainability methods, prompt hygiene, decision provenance, sandboxing, and ModelOps pipelines, to AMAS environments. We critically analyze their limitations and adaptation needs, providing a foundation for re-engineering these techniques to fit the demands of interacting LLM agents (Secs. 79).
  • Building on the identified research gaps, we outline a research agenda that prioritizes scalable, verifiable, and regulation-compliant Agentic AI (Sec. 11).
2 Related Work
Table 1: Comparison of Related Surveys on LLM-Based Multi-Agent Systems and TRiSM Aspects. Existing work offers limited coverage of trust, risk, and security in Agentic AI.
[…]
Guo et al. (2024) [27] ✓ (simulated env’ts) ~ (research challenges)
Chen et al. (2025) [28] ~ (task/simulation focus) ~ (future research areas)
Yan et al. (2025) [29] ✓ (mentions diverse) ~ (future work directions)
Tran et al. (2025) [30] ✓ (networks, QA, etc.) ~ (open challenges)
Lin et al. (2025) [31] ~ (creative tasks) ~ (roadmap for research)
Fang et al. (2025) [32] ~ ~ ✓ (health, finance cited) ~ (technical “next steps”)
Xi et al. (2025) [33] ✓ (single-/multi-agent,
human-AI coop)
~ (open problems)
Luo et al. (2025) [34] ~ ~ ✓ (science, games, etc.) ~ (open challenges)
Zou et al. (2025) [35] ~ ✓ (e.g., finance,
healthcare)
~ (challenges & opp.)
Wang et al. (2025) [36] ~ ~ ~ (healthcare domain) ~ (future opportunities)
This Survey (2025) ✓ (high-stakes domains) ✓ (tailored guidance)
Legend: ✓= explicitly addressed; ~= partially/indirectly addressed; ✗= not addressed.
[…]
We present the related work in this section.

### v4

@@ Abstract
Agentic AI systems, built upon large language models (LLMs) and deployed in multi-agent configurations, are redefining intelligence, autonomy, collaboration, and decision-making across enterprise and societal domains. This review presents a structured analysis of Trust, Risk, and Security Management (TRiSM) in the context of LLM-based Agentic Multi-Agent Systems (AMAS). We begin by examining the conceptual foundations of Agentic AI and highlight its architectural distinctions from traditional AI agents. We then adapt and extend the AI TRiSM framework for Agentic AI, structured around key pillars: Explainability, ModelOps, Security, Privacy and their Lifecycle Governance, each contextualized to the challenges of AMAS. A risk taxonomy is proposed to capture the unique threats and vulnerabilities of Agentic AI, ranging from coordination failures to prompt-based adversarial manipulation. To support practical assessment in Agentic AI works, we introduce two novel metrics: the Component Synergy Score (CSS), which quantifies the quality of inter-agent collaboration, and the Tool Utilization Efficacy (TUE), which evaluates the efficiency of tool use within agent workflows. We further discuss strategies for improving explainability in Agentic AI, as well as approaches to enhancing security and privacy through encryption, adversarial robustness, and regulatory compliance. The review concludes with a research roadmap for the responsible development and deployment of Agentic AI, highlighting key directions to align emerging systems with TRiSM principles-ensuring safety, transparency, and accountability in their operation.
@@ 1 Introduction
AI governance has moved from aspiration to obligation. The EU Artificial Intelligence Act [1] entered into force on 1 August 2024 with a staged application (e.g., prohibitions in early 2025 and obligations for general-purpose AI from 2 August 2025  [^3] EU AI Act); meanwhile, organizations are adopting management frameworks such as ISO/IEC 42001:2023 [2] and the NIST AI Risk Management Framework (AI RMF 1.0) [3] to operationalize risk controls. In parallel, enterprise use of AI continues to expand: recent global surveys [4, 5, 6] report that a substantial majority of organizations use AI in at least one business function. These trends sharpen demand for system-level trust, risk, and security management tailored to agentic systems.
Figure 1: TRiSM taxonomy for Agentic AI (AMAS) presented in this review. We cover: (i) Fundamentals of Agentic AI (Agentic vs. traditional agents; LLM-based architectures), (ii) AI Trust, Risk, and Security Management (TRiSM) pillars: Explainability, ModelOps, Security, Privacy, and Lifecycle Governance, (iii) Trust management via explainability, (iv) Threats and risks, and (v) Evaluation metrics, including human-centered and composite measures.
An AI agent is a computational entity that perceives its environment and acts to achieve goals [7]. What were once task-specific and largely deterministic programs have rapidly evolved into LLM-powered agentic systems with planning, tool use, and persistent memory [8]. In this paper, we refer to such LLM-based, coordinating agent systems as Agentic multi-agent systems (AMAS). These systems orchestrate multiple specialized agents to address long-horizon tasks, coordinate functions and roles, and adapt workflows through interactions with tools, users, and other agents. The shift from single-agent or rule-based pipelines to AMAS unlocks new capabilities but complicates coordination, control, transparency, and risk management across heterogeneous components and contexts.
Why TRiSM for Agentic AI?
LLM-based AMAS exhibit emergent and often opaque behaviors, making them susceptible to cascading errors, biased decisions, and unintended interactions [9]. These risks are amplified when agents plan, coordinate, and call external tools. Beyond traditional ML concerns (safety, fairness, and interpretability), Agentic AI must contend with threats from tool integrations, prompt-level attacks, memory poisoning, impersonation, and privacy leakage in collaborative settings [10, 11]. To address these system-level risks, AI Trust, Risk, and Security Management (TRiSM) frameworks emphasize governance, explainability, security, privacy, and lifecycle controls [12, 13]. We argue that a TRiSM perspective is necessary to make AMAS deployable in sensitive domains.
Scope and Objectives
Our focus is LLM-based AMAS that demonstrate autonomous planning, tool use, memory retention, and emergent reasoning, with or without human oversight. [^4] Throughout, AMAS denotes multi-agent systems built on LLMs. We synthesize recent work on AMAS and AI TRiSM to provide a coherent, actionable reference for researchers and practitioners. Specifically, we map TRiSM pillars :Explainability, ModelOps, Security, Privacy, and their Lifecycle Governance, to AMAS workflows and failure modes. We also surface gaps where current ML governance and model-centric controls do not readily extend to multi-agent settings.
Necessity of this Survey
Much of the Agentic AI literature prioritizes agent modeling, planning, and collaboration, while comparatively under-addressing adversarial robustness, lifecycle governance, decision provenance, and system-level explainability. As deployments expand in healthcare, finance, science, and public services, the absence of an integrated TRiSM perspective exposes stakeholders to opaque decision pathways and unmanaged risks. To our knowledge, few surveys examine LLM-based multi-agent systems explicitly through the lens of trust, risk, security, and governance; Table 1 situates our contribution relative to prior reviews.
Table 1: Comparison with related surveys.
Survey Threats Lifecycle Gov. Explainability TRiSM Integration LLM-Specific Applications Actionable Guidance
Guo et al. (2024) [14] ~
Chen et al. (2025) [15] ~ ~
Yan et al. (2025) [16] ~
Tran et al. (2025) [17] ~
Lin et al. (2025) [18] ~ ~
Fang et al. (2025) [11] ~ ~ ~
Xi et al. (2025) [19] ~
Luo et al. (2025) [20] ~ ~ ~
Zou et al. (2025) [21] ~ ~
Wang et al. (2025) [22] ~ ~ ~ ~
This Survey (2025)
Legend: ✓= explicitly addressed; ~= partially addressed; ✗= not addressed.
Contributions
We make the following contributions:
  • Unified TRiSM framework for Agentic AI. We propose and present a system-level Trust, Risk and Security Management (TRiSM) framework tailored to large-language-model-based multi-agent systems (Section 5). The framework integrates explainability, ModelOps, security, privacy and lifecycle governance, contextualising each pillar to AMAS workflows and failure modes.
  • Risk taxonomy for AMAS. We synthesise a taxonomy of threats unique to LLM-driven multi-agent systems, such as prompt injection, memory poisoning, collusive failure, emergent misbehaviour and tool-use abuse; and map these risks to corresponding controls (Section 4).
  • Evaluation template and new metrics. We propose a holistic evaluation template that covers trustworthiness, explainability, user-centred performance and inter-agent coordination. To operationalise this template, we introduce the Component Synergy Score (CSS) for measuring collaboration quality across agents and Tool Utilisation Efficacy (TUE) for assessing the correctness and efficiency of tool calls (Section 6).
  • Technique mapping and limitations. We survey and map existing explainability methods, prompt-hygiene practices, decision-provenance tools, sandboxing strategies and ModelOps pipelines to the AMAS setting. This mapping reveals where existing approaches can be adapted and where significant gaps remain (Section 8).
Building on the above contributions, we outline a research roadmap for developing scalable, verifiable and regulation-aligned agentic systems. Key directions include adversarial robustness, governance protocols, standardised benchmarks for trustworthiness and coordination, and cross-disciplinary collaboration.
2 Background, Related Work and Literature Review Methodology
2.1 AI Agents vs. Agentic AI
@@ 2 Background, Related Work and Literature Review Methodology
An AI agent, typically defined as a computational entity that perceives its environment and takes actions to achieve goals [7], has undergone a rapid transformation in recent years. While early agents were task-specific and deterministic, modern systems have evolved substantially, increasing operational autonomy on benchmark tasks [23]. These agentic systems integrate large language models (LLMs), external tools, and persistent memory to support complex planning and coordinated decision-making [8]. This evolution presents new opportunities but also raises challenges in coordination, control, transparency, and risk assessment across heterogeneous agents. As the shift toward agentic architectures accelerates, long-standing concerns around safety, fairness, interpretability, and systemic bias resurface [11], especially in high-stakes domains such as healthcare, law, and public services, where decisions can carry profound societal consequences.
@@
Autonomous software agents date back to the 1990s [24]; however, the advent of LLM-driven multi-agent ecosystems since 2023 [9] has made “Agentic AI” a distinct and rapidly expanding research focus [8]. Traditional agents primarily automated limited tasks—such as information retrieval [25], data summarization [26], and dialogue response [27]—often via single-step logic [28] or scripted rules. These approaches lacked the deep reasoning, adaptability, and persistence required for complex, multi-step problem-solving.
@@
In contrast, agentic systems coordinate collaborative agents with specialized roles (e.g., planner, coder, analyst) enabled by LLMs and tool use [29]. They dynamically decompose tasks, share context, and pursue high-level goals over long horizons [30]. This shift reflects not only a change in technology but also a fundamental increase in complexity and autonomy, marking the emergence of machine collectives capable of exhibiting emergent, decentralized behavior.
@@ 2.2 Trust, Risk, and Security Challenges in Agentic AI
The rise of Agentic AI introduces challenges that extend beyond those of traditional ML systems. Unlike deterministic agents [31], LLM-based multi-agent systems (AMAS) exhibit emergent and often opaque decision-making, increasing the risk of cascading errors, biases, and unintended behaviors. These characteristics necessitate robust approaches to building trust and ensuring trustworthiness in high-stakes domains such as healthcare, finance, and law. Below, we discuss several key challenges:
@@
    Trust and trustworthiness. In Agentic AI, trust refers to a user’s willingness to rely on an AI system, while trustworthiness denotes whether the system consistently behaves in a safe, fair, and predictable manner. According to ISO [32], trustworthiness is a quality characteristic of systems and is understood as the ‘ability to meet stakeholders’ expectations in a verifiable way’. This distinction is critical in high-stakes settings, where undue reliance on an untrustworthy system can lead to severe consequences, such as misdiagnoses in healthcare or biased rulings in legal applications. The stochastic nature of LLMs [33] further complicates trustworthiness: randomness in output generation, while sometimes promoting creativity, introduces inconsistency and limits repeatability.
@@
    Security risks in multi-agent systems. Risk surfaces expand as autonomous agents gain access to tools, external APIs, and persistent memory, amplifying the potential for privacy breaches, adversarial misuse, or regulatory violations in the absence of robust human oversight [10]. For example, a multi-agent supply-chain optimizer may coordinate procurement and logistics agents yet inadvertently leak sensitive data or violate compliance protocols (e.g., [34], HIPAA [35]) if safeguards are inadequate.
@@
    Prompt injection, spoofing, and impersonation. AMAS are susceptible to security threats that exploit language-based interfaces and cooperative behaviors. One prominent threat is prompt injection [36], where attackers craft inputs with hidden or malicious instructions to manipulate agent behavior. Recent studies [37] describe “prompt infection”, in which malicious prompts propagate from one agent to another, akin to a computer virus spreading across networks [38]. Corrupted outputs can cascade, leading to data leaks, fraudulent transactions, misinformation, or coordinated misbehavior within an agent society.
    Another critical vulnerability involves spoofing and impersonation [39]. Spoofing refers to falsifying identity or credentials (e.g., using a false name or address) to deceive a system, whereas impersonation involves pretending to be a specific agent or user to gain trust, access, or privileges. In AMAS, where agents rely on credentials or tokens for authentication during coordination, an adversary who steals an API key or mimics a trusted peer can issue unauthorized commands or access sensitive information.
@@
    Privacy risks in multi-agent LLM systems. LLM-based agents often process user data, proprietary business information, and other sensitive inputs to fulfill their tasks [40]. In a multi-agent context, this challenge is amplified by inter-agent information sharing (e.g., via a shared memory store or message passing), which complicates data minimization, access control, and purpose limitation.
@@
Traditional evaluation and safety frameworks, designed for deterministic, single-agent pipelines; are inadequate for the autonomy and control properties of AMAS. To address these gaps, AI Trust, Risk, and Security Management (AI TRiSM) [12, 13] proposes lifecycle controls spanning explainability, secure model/tool orchestration, and privacy management. Rather than focusing solely on technical performance, TRiSM integrates governance practices to strengthen reliability and robustness, protect data, and align system behavior with ethical and regulatory expectations.
2.3 Survey Landscape